[Brahms-l] Shift report 20050314 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Mar 14 2005 - 16:01:37 EST
Shifter leader was Hiro
Shifters were 


1.  Beam for most of shift
2.  
    FV adjusted timing for trig4.  see elog.
    trig 5 was included.
    There were several different settings in this shift. See elog for detail.
    MRS was moved to 45 deg from run 13783.  
    FS was at 6 deg until 13783 1/4A.
    FS was at 8 deg from Run 13784 to 13789 1/4A.
    FS was at 6 deg from 13790 1/4B.
3.  Not detected.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Mar 14 16:01:49 2005

This archive was generated by hypermail 2.1.8 : Mon Mar 14 2005 - 16:01:56 EST